# TLDP: An Algorithm of Local Differential Privacy for Tensors

### v1

@@ III-B Randomized Response
Randomized response is a statistical survey technique introduced in the 1960s for gathering data on sensitive topics while allowing respondents to maintain confidentiality. Surveys employing this method enable the accurate calculation of precise population statistics while ensuring individual privacy. Today, it is the dominant perturbation mechanism used in local differential privacy protection. For binary-valued problems, the randomized response method retains the original data with probability $p$ and flips it with probability $1-p$. In most cases, $p>\frac{1}{2}$, $i.e.p>1-p$, so for any $v_{1}\neq v_{2}$, we have
@@
Hence it satisfies $\ln(\frac{p}{1-p})$-LDP.
IV Local Differential Privacy of Tensors (TLDP)
@@
In LDP, users are unaware of each other’s records, meaning the concept of global sensitivity, which measures the maximum change in query results between two adjacent datasets in differential privacy, does not apply. In this paper, we use $\Delta$ to denote the range of variation of the raw data. It is important to note that $\Delta$ differs from global sensitivity: the former is only related to the users’ initial data, while global sensitivity pertains to query functions and is affected by the size of tensors.
Comparing Definition 1 with Definition 2, it is evident that Definition 1 is a special case of Definition 2, where the tensor order is one ($i.e.,N=1,I_{1}=1$). However, the randomized response applied to clients in [12] does not adhere to the $\epsilon$-LDP definition presented here. In this approach, the central server receives the original, unperturbed data from a random user with a certain probability. From the perspective of definition, when only one tensor between $\mathcal{X}$ and $\mathcal{X}^{\prime}$ is perturbed, the privacy budget $\epsilon$ becomes very large or even approaches infinity. This suggests that the method may fail to provide efficient privacy protection with a certain probability.
IV-A Local Differential Privacy for Tensors (TLDP)
We adopt the widely used randomized response technique to implement LDP on tensors. This approach allows for the aggregation of collective data while safeguarding individual privacy, eliminating the need for a trusted central server by introducing random flips of client-side information. Originally tailored for scalar data with binary outcomes, the randomized response technique has been extended to multi-valued scenarios by encoding scalars into higher-dimensional vectors or adjusting the flipping probabilities.
@@ IV-A Local Differential Privacy for Tensors (TLDP)
However, applying these methods to tensor data presents unique challenges. On the one hand, encoding each tensor element individually would demand substantial storage resources, exponentially increasing algorithmic complexity. On the other hand, given our research’s requirement to leverage information from each tensor rather than their aggregation, methods that randomly select tensor elements for perturbation could severely reduce the utility of the original data.
Previous studies on tensor privacy protection have not fully resolved the issue of implementing LDP for tensor data. For example, [11] achieves differential privacy by appending a noise tensor of equal magnitude to the data tensor. However, LDP demands a higher level of privacy than traditional differential privacy, necessitating more noise to achieve comparable privacy guarantees, which can significantly degrade data utility due to extensive perturbation. [12] applies differential privacy to high-order, high-dimensional sparse tensors within the IoT transmission context through individual randomized responses. But when the central server is required to receive tensor information directly from clients, this approach risks disclosing the user’s original information with a certain probability, thereby undermining the purpose of privacy protection.
@@
In light of these challenges, there is a clear need for more effective methods that balance privacy protection and the utility of tensor data. Building upon the aforementioned concepts, we propose a novel approach for achieving tensor differential privacy: selectively perturbing specific positions with noise determined via the randomized response technique. In this method, each tensor element remains its original value with a probability of $p$ and introduces calibrated noise (either Laplace or Gaussian) with a probability of $1-p$. Our method ensures privacy protection for each user and demonstrates that compared to directly adding a noise tensor, the inherent randomness of randomized response allows for meeting privacy budgets without requiring excessive noise, thereby maintaining a balance between privacy preservation and data utility.
@@
Let $\mathcal{X}$ and $\mathcal{X}^{\prime}$ be any two tensors of the same size, and $\mathcal{X}^{\ast}$ be any output. It can be observed that the probability of the tensor remaining unperturbed is the highest. We assume that in the unperturbed tensor $\mathcal{X}$, there are a total of t components that preserve the original data, as in step 8 in Algorithm 1.
@@
    $\frac{Pr[\mathcal{M}_{Lap}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Lap}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$
@@
    $N(x|\sigma^{2})=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}},$
@@
    $\frac{Pr[\mathcal{M}_{Gau}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Gau}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$
@@
    $\frac{Pr[\mathcal{M}_{Lap}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Lap}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ $\leq\frac{(\frac{1}{2b_{1}})^{I}}{(\frac{1}{2b_{1}}\cdot e^{-\frac{\Delta}{b_{1}}})^{I}}$
@@
    $\frac{Pr[\mathcal{M}_{Gau}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Gau}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ $\leq\frac{(\frac{1}{\sigma_{1}\sqrt{2\pi}})^{I}}{(\frac{1}{\sigma_{1}\sqrt{2\pi}}\cdot e^{-\frac{\Delta^{2}}{2\sigma_{1}^{2}}})^{I}}$
@@
    $\frac{\sigma^{2}}{\sigma^{2}_{1}}=\frac{1}{I}$
@@ IV-B Weighted TLDP
Let $\mathcal{X}$ and $\mathcal{X}^{\prime}$ be any two tensors of the same size, and $\mathcal{X}^{\ast}$ be any output. It can be observed that the probability of the tensor remaining unperturbed is the highest. We assume that in the unperturbed tensor $\mathcal{X}$, there are a total of $t$ components that preserve the original data, as in step 9 in Algorithm 2.
@@
    $\frac{Pr[\mathcal{M}_{Lap}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Lap}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$
@@
    $N(x|\sigma^{2})=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}},$
@@
    $\frac{Pr[\mathcal{M}_{Gau}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Gau}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$
@@ V Evaluations
To provide a fair and comprehensive assessment of the mechanisms we propose, we have conducted an extensive series of experiments across various models and datasets. The results are then compared with those of several existing mechanisms, offering detailed insights into their relative performance.
@@ V-A1 Datasets and Models
In this study, we have carefully selected representative learning tasks from diverse domains, including computer vision, text mining, and data mining, where the datasets are inherently sensitive. Given the confidentiality and proprietary nature of the data, it is more secure for data owners if this information is perturbed before being transmitted to the central server.
@@
For the computer vision task, we select MNIST [37], CIFAR-10 [38], and SVHN [39] as the datasets. The MobileNet model, utilizing the pre-trained version available in torchvision, is employed. The model is trained with the Adam optimizer, a learning rate of 0.001, cross-entropy loss, a batch size of 64, and 10 epochs. The MNIST dataset consists of 60,000 training samples and 10,000 test samples of 28$\times$28 pixel grayscale images of handwritten digits. To adapt the dataset for use with the MobileNet model, each single-channel image is replicated across three channels with identical content. Both CIFAR-10 and SVHN contain 50,000 training samples and 10,000 test samples, consisting of 10 classes of RGB images. CIFAR-10 features images of common objects from the real world, while SVHN includes door number images collected from Google Street View.
@@
In the context of data mining, we employ the DarkNet dataset [40], which is designed to detect and characterize VPN and Tor applications. It combines two public datasets from the Canadian Institute for Cybersecurity (CIC), namely ISCXTor2016 and ISCXVPN2016. The Tor data is divided into 200 segments, which are then interspersed with non-Tor data. The final 20% of the combined dataset is used for testing, with the remaining 80% used for training, resulting in 113,184 training samples and 28,297 test samples. For this task, a Fully Connected Network (FCN) model is employed. The model consists of four hidden layers, with the first three having a hidden size of 256 and the last layer having a hidden size of 128. A dropout rate of 0.3 is applied. The model is trained using the Adam optimizer, with a learning rate of 0.001, cross-entropy loss, a batch size of 64, and 25 epochs.
@@
For text mining, we work with the IMDB dataset [41], which consists of 50,000 movie reviews used for binary classification. Each review is tokenized using the spaCy tokenizer, and each token is converted to its corresponding vocabulary index using pre-trained GloVe embeddings, with each review truncated or padded to a fixed length of 500 tokens. Both the training and test sets consist of 25,000 samples each, and the data is directly imported from torchtext. A BiLSTM model with two hidden layers, each with a hidden size of 256, is employed, with a dropout rate of 0.3. The model is trained using the Adam optimizer, a learning rate of 0.001, cross-entropy loss, a batch size of 64, and 50 epochs.
V-A2 Baselines
@@
We compare our method with other differential privacy mechanisms designed for high-dimensional data. The baselines include the Laplace mechanism (Laplace), Gaussian mechanism (Gaussian), Matrix-Variate Gaussian (MVG) [20], and Tensor Variate Gaussian (TVG) [11]. Unlike previous studies where noise designed for scalar-valued queries is directly applied to tensors, we have mathematically proven that the Laplace and Gaussian noise we add satisfies LDP for tensors (Definition 2), as detailed in Theorem 2. TVG, on the other hand, is a $(\epsilon,\delta)$-differential privacy mechanism for tensor-valued queries that use directional noise. It is important to note that the TVG mechanism we use satisfies $(\epsilon,\delta)$-DP, rather than LDP as described above. Nonetheless, we include it in our comparison to prove the effectiveness of our approach. Specifically, we use Independent Directional Noise (IDN) [11] as the perturbation method. In all experiments, the utility subspace $W$ is set to the identity matrix $E$, as no prior knowledge is available.
Although the k-RR technique proposed by Kairouz et al. [13] demonstrates superior performance in multi-valued perturbation under single-value frequency publication, it heavily relies on the central server’s prior knowledge of the users’ variable values. To address the challenge, they introduced hash functions and grouping operations based on the k-RR algorithm, resulting in the $O$-RR method. This adaptation significantly enhances the efficiency of frequency estimation. However, in our application scenario, the server may not have prior knowledge of user variable values, making the k-RR algorithm unsuitable. Our experimental goal is to perform federated learning with perturbed data, and due to the irreversible nature of hash functions, the $O$-RR method is unsuitable in this application scenario. As a result, both k-RR and $O$-RR methods are excluded from our experimental comparisons.
V-A3 Metrics
In our experiments, we assess the performance of classification models using several key metrics: accuracy, weighted recall, weighted precision, and weighted F1-score.
Accuracy measures the model’s overall correctness, defined as the ratio of correctly predicted instances to the total number of instances. Weighted recall, on the other hand, calculates the average recall for each class but assigns weights proportional to the number of samples in each class. This ensures that classes with more samples contribute more significantly to the final score, reflecting recall across all classes while emphasizing those with larger sample sizes. This approach is particularly useful for imbalanced datasets, where certain classes may have substantially fewer samples than others. Similarly, weighted precision computes the average precision for each class, weighted by the number of samples in each class. Like weighted recall, it addresses class imbalance by prioritizing classes with more samples. The weighted F1-score, the harmonic mean of weighted precision and weighted recall, provides a balanced evaluation of both metrics, with weights determined by the sample size of each class. This ensures a fair assessment of the model’s performance across all classes, regardless of their size.
By employing these metrics, we aim to deliver a comprehensive evaluation of the model’s performance, thoroughly examining its capabilities across various dimensions. For clarity, throughout this paper, the terms recall, precision, and F1-score refer to their weighted versions unless otherwise specified.
V-A4 Privacy-Preserving Targets
In diverse real-world application scenarios, different tasks may have distinct privacy preservation objectives. We categorize these objectives into three primary types: original testing data, training features, and model gradients. For the first type, users simply transmit the raw data to the server, relying on the server’s capabilities to process and analyze the data without the need for additional feature extraction or model training on the client side. The second type focuses on features extracted from the original data. Feature extraction is an effective method for dimensionality reduction, which not only reduces communication costs but also uncovers more insightful underlying variables, thereby enhancing the understanding of the data. The third type emphasizes the protection of model gradients, which are critical for updating model parameters during training. In particular, we utilize the noisy Stochastic Gradient Descent (SGD) [21] as our optimization algorithm, where users compute their own gradients before transmitting them to the central server.
In most tensor-related simulation experiments, tensors are typically subjected to a ”flattening” process. However, in previous work, regardless of whether the user’s information is scalar or tensor, the noise introduced to the raw data is deeply affected by the sensitivity of the query function, one key factor of which is the size of the tensor. Essentially, the larger the specifications of the data tensors, the more substantial the noise tends to be, potentially leading to a decrease in the accuracy of the results. For machine learning tasks outlined in [11], deep learning efforts discussed in [21] and the efficient estimation of high-dimensional joint distributions as covered in [42], the exact details of the original data are not essential. As a result, techniques such as convolution can be employed to reduce tensor dimensions, effectively mitigating the problem of excessive noise. So in types 2 and 3, features and gradients are transmitted and trained as information between users and the central server. However, since dimensionality reduction invariably involves a loss of the original message, resulting in a decrease in model accuracy, it becomes imperative in certain scenarios to learn from complete data. Therefore, in Type I, disturbed tensors are sent to the central server without any processing. We present the implementation details categorized by type as follows.

### v2

@@ III-B Randomized Response
Randomized response is a statistical technique introduced in the 1960s by [3] for gathering data on sensitive topics while allowing respondents to maintain confidentiality. Surveys employing this method enable the accurate calculation of precise population statistics while ensuring individual privacy. Today, it is the dominant perturbation mechanism used in local differential privacy protection of scalars. For binary-valued problems, the randomized response method retains the original data with probability $p$ and flips it with probability $1-p$. In most cases, $p>\frac{1}{2}$, i.e., $p>1-p$, so for any $v_{1}\neq v_{2}$, we have
@@
It satisfies $\ln(\frac{p}{1-p})$-LDP.
IV Local Differential Privacy of Tensors
@@
In Local Differential Privacy (LDP), users are unaware of each other’s records. As a result, the concept of global sensitivity, which quantifies the maximum change in query results between two adjacent datasets in differential privacy, does not apply. In this paper, we use $\Delta$ to denote the range of variation of the raw data. It is important to note that $\Delta$ differs from global sensitivity: the former is only related to the users’ initial data, while global sensitivity pertains to query functions and is affected by the size of tensors.
Comparing Definition 1 with Definition 2, it is evident that Definition 1 is a special case of Definition 2, where the tensor order is one (i.e., $N=1,I_{1}=1$).
IV-A TLDP
We introduce TLDP to protect the privacy of a user’s tensor data in distributed systems before transmitting it to untrusted parties. This approach enables the aggregation of collective data while preserving individual privacy, eliminating the need for a trusted central server.
@@ IV-A TLDP
Originally designed for scalar data with binary outcomes, the randomized response technique has been extended to multi-valued scenarios by encoding scalars into higher-dimensional vectors or adjusting flipping probabilities. However, applying these methods to tensor data presents unique challenges. Encoding each tensor element individually would require substantial storage resources, exponentially increasing algorithmic complexity. Conversely, given our research’s need to retain information from each tensor rather than relying solely on aggregation, methods that randomly perturb selected tensor elements could significantly reduce the utility of the original data.
@@
In light of these challenges, there is a clear need for more effective methods that balance privacy protection and the utility of tensor data. Building upon the aforementioned concepts, we propose a novel approach named TLDP for achieving tensor differential privacy by selectively perturbing specific positions with noise determined via randomized response. In this method, each tensor element remains its original value with a probability of $p$ and introduces calibrated noise (either Laplace or Gaussian) with a probability of $1-p$. Our method ensures privacy protection for each user and demonstrates that compared to directly adding a noise tensor, the inherent randomness of randomized response allows for meeting privacy budgets without requiring excessive noise, thereby maintaining a balance between privacy preservation and data utility.
@@
Let $\mathcal{X}$ and $\mathcal{X}^{\prime}$ be any two tensors of the same size from the same user, and $\mathcal{X}^{\ast}$ be any output. It can be observed that the probability of the tensor remaining unperturbed is the highest. We assume that in the unperturbed tensor $\mathcal{X}$, there are a total of t components that preserve the original data, because the corresponding random numbers are smaller than p.
@@
    $\frac{Pr[\mathcal{M}_{Lap}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Lap}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ (2)
@@
    $N(x|\sigma^{2})=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}},$ (3)
@@
    $\frac{Pr[\mathcal{M}_{Gau}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Gau}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ (4)
@@
    $\frac{Pr[\mathcal{M}_{Lap}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Lap}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ $\leq\frac{(\frac{1}{2b_{1}})^{I}}{(\frac{1}{2b_{1}}\cdot e^{-\frac{\Delta}{b_{1}}})^{I}}$ (5)
@@
    $\frac{Pr[\mathcal{M}_{Gau}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Gau}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ $\leq\frac{(\frac{1}{\sigma_{1}\sqrt{2\pi}})^{I}}{(\frac{1}{\sigma_{1}\sqrt{2\pi}}\cdot e^{-\frac{\Delta^{2}}{2\sigma_{1}^{2}}})^{I}}$ (7)
@@
    $\frac{\sigma^{2}}{\sigma^{2}_{1}}=\frac{1}{I}.$ (8)
@@ IV-B Weighted TLDP
Let $\mathcal{X}$ and $\mathcal{X}^{\prime}$ be any two tensors of the same size, and $\mathcal{X}^{\ast}$ be any output. It can be observed that the probability of the tensor remaining unperturbed is the highest. We assume that in the unperturbed tensor $\mathcal{X}$, there are a total of t components that preserve the original data, because the corresponding random numbers are smaller than p.
@@
    $\frac{Pr[\mathcal{M}_{Lap}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Lap}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ (10)
@@
    $N(x|\sigma^{2})=\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{x^{2}}{2\sigma^{2}}},$ (11)
@@
    $\frac{Pr[\mathcal{M}_{Gau}(\mathcal{X})=\mathcal{X}^{\ast}]}{Pr[\mathcal{M}_{Gau}(\mathcal{X}^{\prime})=\mathcal{X}^{\ast}]}$ (12)
@@ V Evaluations
To provide a fair and comprehensive assessment of the proposed mechanisms, we conduct extensive experiments across various models and datasets. The results are compared with several existing mechanisms, offering detailed insights into their relative performance.
@@ V-A1 Datasets and Tasks
In this study, we carefully select representative learning tasks from diverse domains, including computer vision, text mining, and data mining, where the datasets are likely to be sensitive. Given the confidentiality and proprietary nature of the data, it is more secure for data owners if this information is perturbed before being transmitted to the untrusted central server. For example, raw data or features extracted from raw data could contain sensitive information (e.g., personal images or house numbers), and this information should be carefully preserved before being transmitted to third parties. Another example is that the model parameters that are trained on sensitive data can also be sensitive, and the deep learning models computed on such data should not reflect it.
@@
For the computer vision task, we select MNIST [36], CIFAR-10 [37], and SVHN [38] as the datasets. MobileNet (pre-trained on ImageNet)  [^1] https://pytorch.org/vision/main/models/generated/torchvision.models.mobilenet_v2.html#torchvision.models.MobileNet_V2_Weights is used for the computer vision task. The model is trained with the Adam optimizer with a learning rate of 0.001. It uses the cross-entropy loss and has a batch size of 64. We train it for 10 epochs. The MNIST dataset consists of 60,000 training samples and 10,000 test samples of 28$\times$28 pixel grayscale images of handwritten digits. CIFAR-10 contains 50,000 training samples and 10,000 test samples and has 10 classes of common objects, e.g., birds, airplanes, and cats. SVHN includes door number images collected from Google Street View, and it has 73,257 training images and 26,032 testing images.
@@
In the context of data mining, we employ the DarkNet dataset [39], which is designed to detect and characterize VPN and Tor applications. It combines two public datasets from the Canadian Institute for Cybersecurity (CIC), namely ISCXTor2016 and ISCXVPN2016. The Tor data is divided into 200 segments, which are then interspersed with non-Tor data. The final 20% of the combined dataset is used for testing, with the remaining 80% used for training, i.e., 113,184 training samples and 28,297 test samples. For this task, a Fully Connected Network (FCN) is employed, which consists of four hidden layers, with the first three having a hidden size of 256 and the last layer having a hidden size of 128. A dropout rate of 0.3 is applied. The model is trained using the Adam optimizer with a learning rate of 0.001. It also uses the cross-entropy loss with a batch size of 64. We train it for 25 epochs.
@@
For the text mining task, we work with the IMDB dataset [40], which consists of 50,000 movie reviews used for binary classification. Each review is tokenized using the spaCy tokenizer, and each token is converted to its corresponding vocabulary index using pre-trained GloVe embeddings, with each review truncated or padded to a fixed length of 500 tokens. Both the training and test sets consist of 25,000 samples each, and the data is directly imported from Torchtext. A BiLSTM model with two hidden layers, each with a hidden size of 256, is employed with a dropout rate of 0.3. The model is trained using the Adam optimizer with a learning rate of 0.001. The cross-entropy loss is utilized for model training with a batch size of 64, and we train the model for 50 epochs.
V-A2 Baselines and Metrics
@@
We compare TLDP with other differential privacy mechanisms designed for high-dimensional data. The baselines contain: the Laplace mechanism (Laplace), Gaussian mechanism (Gaussian), Matrix-Variate Gaussian (MVG) [16], and Independent Directional Noise (IDN) [17]. We implement TLDP-L, TLDP-G, TLDP-L-w, and TLDP-G-w as solutions for applying local differential privacy on private data (see Section IV), where TLDP-L and TLDP-G denote TLDP that uses the Laplace and Gaussian noise while TLDP-L-w and TLDP-G-w represent TLDP when using Laplace or Gaussian noise while applying a weight matrix for improving utility. Unlike previous studies where noise designed for scalar-valued queries is directly applied to tensors, we mathematically prove that the Laplace and Gaussian noise we add satisfies LDP for tensors (Definition 2), as detailed in Theorem 2. IDN, on the other hand, is a $(\epsilon,\delta)$-differential privacy mechanism for tensor-valued queries. In all experiments, the utility subspace $W$ is set to the identity matrix $E$.
In our experiments, we assess the performance of classification models using a weighted F1-score (i.e., F1-score in the following text) for fair comparison over various tasks, including tasks with imbalanced distributions.
The F1-score is calculated as follows:
$P_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}},\$ (13)
$R_{i}=\frac{TP_{i}}{TP_{i}+FN_{i}},\$ (14)
where $TP$ (True Positives) means the number of instances where the model correctly predicts the positive class. $FP$ (False Positives) is the number of instances where the model incorrectly predicts the positive class (i.e., it predicts positive when the actual class is negative). $FN$ (False Negatives) is the number of instances where the model incorrectly predicts the negative class (i.e., it predicts negative when the actual class is positive).
$F1_{i}=2\times\frac{P_{i}\times R_{i}}{P_{i}+R_{i}},\$ (15)
[…]
where $w_{i}=\frac{N_{c}}{N}$. Here, $N_{c}$ and $N$ denote the number of samples of class $c$ and the total number of samples from all the classes, respectively.
V-A3 Privacy-Preserving Targets
In our experiments, we simulate the distributed computation system using three users and one central server. The users compute their local models using their private dataset and transmit the trained model parameters to the central server. The central server aggregates the received model parameters, computes an average of them, and sends it back to all the users. We denote the models trained with each user’s dataset as local models and the aggregated model on the central server as the global model. Here, we consider three typical tasks in the distributed computation system, i.e., original data, training features, and model parameters. For the first type, users simply transmit their private data to the server, leveraging the enormous computing resources of the server to process and analyze the data. The second type focuses on features extracted from private datasets. This scenario frequently arises in distributed model training across multiple parties, where users process raw data to extract features and transmit the extracted feature embeddings instead of raw data to the central server for model training to avoid privacy leakage. The query mechanism outputs intermediate data/features with differential privacy guarantees, which are then utilized for subsequent training processes on the server side. The third type emphasizes the protection of model parameters. Sharing model parameters is very common in distributed systems, especially when federated learning [41] is evolved for model computation without the need for direct data or feature sharing between data owners and a parameter server. However, research [42] has found that attackers could easily infer the original data used to train the model from intercepted model parameters. Given that the trained model parameters are sensitive, the query function returns a differentially private version of them before transmitting. We present the implementation details categorized by type as follows.